POlicies Current Attempt in Progress Ivanhoe Company bottles and distributes B-Lite, a diet soft drink. The beverage is soldPrepare a CVP income statement for 2020 based on managements estimates. IVANHOE COMPANY CVP Income Statement (Estimated) For$Calculate variable cost per bottle. (Round variable cost per bottle to 3 decimal places, e.g. 0.251.) Variable cost per bottlCompute the break-even point in (1) units and (2) dollars. (Round answers to 0 decimal places, eg. 1,225.) (1) Compute the brCompute the contribution margin ratio and the margin of safety ratio. (Round variable cost per bottle to 3 decimal places, egDetermine the sales dollars required to earn net income of $210,000. (Round answer to 0 decimal places, e.g. 1,225.) Required
Answer #1

1. IVANHOE company CVP Income statement (Estimated) for the year ending 31st December 2020

Sales $19,50,000
Variable costs
COGS $11,70,000
Selling exp 171,000
Adm exp 24,000
Total VC (13,65,000)
Contribution margin 585,000
Fixed costs
COGS 230,000
Selling 59,000
Adm 120,500
Total FC (409,500)
Net income 175,500

2. No. Of units sold = sales / SP per unit = 19,50,000/0.5 = 3900000

VC total 13,65,000 hence VC per unit = 13,65,000/3900000 = 0.35/unit

3.BEP(units) = Fixed costs /contribution margin per bottle

=409500/(0.5-0.35) = 27,30,000 units

BEP($) = FC / Contribution margin ratio

= 409,500/30% = 13,65,000

4. Contribution margin ratio = 585,000/19,50,000 =30%

MOS ratio = (1950,000-13,65,000)/19,50,000 =30%

5. Sales required to earn net income of $240,852

= $409,500 +$240,852/30% = $21,67,840